OpenStudy (anonymous):

Which of the following is not a step used when graphing the equation y = 1/4−x + 8 using the slope-intercept method? Plot the x-intercept (32, 0). Plot the y-intercept (0, 8). Draw a line through the two points. Count down 1 and right 4 from the y-intercept.

OpenStudy (inkyvoyd):

Hey, you've asked a few of these questions. Want to try to work it out on your own, and tell us what you don't get?

OpenStudy (anonymous):

i just dont get how to graph them

OpenStudy (inkyvoyd):

Can you tell me which points are the y and x intercepts?

OpenStudy (anonymous):

y=(0,0) and x=(-1,4)

OpenStudy (anonymous):

is that right?

OpenStudy (inkyvoyd):

Is the equation y = 1/4−x + 8 or y = -(1/4)x + 8?

OpenStudy (anonymous):

the first one

OpenStudy (inkyvoyd):

Ok, first simplify the equation.

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

just like give me an example

OpenStudy (inkyvoyd):

y = 1/4−x + 8 y=-x+32/4+1/4 combine the constants 8 and 1/4

OpenStudy (anonymous):

oh so you multiply 4 & 8 to get the product of 32

OpenStudy (inkyvoyd):

No..

OpenStudy (inkyvoyd):

I rewrote 8 as 32/4

OpenStudy (anonymous):

..... :(

OpenStudy (anonymous):

oh

OpenStudy (inkyvoyd):

Look, I think you have the problem wrong. \(y = \frac{1}{4}−x + 8\) or \(y = -\frac{1}{4}x + 8\)

OpenStudy (anonymous):

#2

OpenStudy (inkyvoyd):

Ok. Then, can you tell me the slope?